OpenStudy (anonymous):

medal and honor Solve. 6(a + 3) = 18 + 6a (Points : 1) no solution all real numbers –3 –1

OpenStudy (juana02):

all are real numbers

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

wuuuut

OpenStudy (juana02):

i solve it gave me all as real numbers

OpenStudy (mathstudent55):

No.

OpenStudy (mathstudent55):

6(a + 3) = 18 + 6a Distribute: 6a + 18 = 18 + 6a Apply commutative property to right side: 6a + 18 = 6a + 18 Subtract 6a and 18 from both sides: 0 = 0 Since 0 = 0 is a true statement, that means this equation is true for all real numbers.

OpenStudy (anonymous):

but that also means no solution

OpenStudy (anonymous):

well at least thats what i learned

OpenStudy (mathstudent55):

@prguan You solved the equation correctly and got 0 = 0. Your only mistake was reaching the wrong conclusion.

OpenStudy (anonymous):

ohhhhhh i get it

OpenStudy (mathstudent55):

You can't have learned that because it's incorrect. What you learned was a similar thing you are now confusing with.

OpenStudy (anonymous):

whoops

OpenStudy (mathstudent55):

There is no solution is the variables are eliminated and the statement is false.

OpenStudy (anonymous):

then indefinitely many solutions

OpenStudy (anonymous):

which also means all real numbers?

OpenStudy (mathstudent55):

For example: 6(a + 3) = 15 + 6a 6a + 18 = 15 + 6a 18 = 15 Since we have a false statement, there is no solution.

OpenStudy (mathstudent55):

Yes. Infinite number of solutions means the same as all real numbers.

OpenStudy (anonymous):

oh ok i thought it indefinitely many solutions but didn't know it as all real numbers